How many liters of an 8% solution of salt should be added to a 25% solution in order to obtain 510 liters of an 12% solution?
I tried many different ways to solve this probelm:
.08x +.25(510-x)=510?
.25x+.08(.12-x)=510?
but none of this come out with the correct answer if you could please help i would greatly appreciate it! Also if you could possibly tell me where i am going wrong? THANK YOU!! In advanced

Answer by chessace(471)

Your 2nd equation is too far gone.
The first one is close, you left out the 12%:
Salt: .08x +.25(510-x)=.12*510
-.17x = -.13*510
x = (13/17)*510 = 13*30 = 390 L